# Adapters for Altering LLM Vocabularies: What Languages Benefit the Most?

### v2

@@ 3.1 Vocabulary Adapter Module
We introduce the vocabulary adapter modules to find parameters of new embeddings that can replace the original embedding without changing the non-embedding part of the original model (Figure 1(a)). For simplicity, we refer to both input and output embeddings (or the LM head) collectively as embeddings. Let $V^{o}$ and $V^{n}$ be the $\bm{o}$riginal and $\bm{n}$ew vocabulary, respectively, and let $\mathcal{T}^{x}:w\rightarrow(t_{1},t_{2},\dots,t_{k})$ be a tokenizer associated with a vocabulary $V^{x}$ where $t_{j}\in V^{x},\forall j=1,\dots,k$. We put vocabulary adapter modules ${\bm{A}}\in\mathbb{R}^{|V^{n}|\times|V^{o}|}$ between the new vocabulary $V^{n}$ and the original embedding ${\bm{E}}^{o}\in\mathbb{R}^{|V^{o}|\times h}$ where $h$ is an embedding dimension, in a manner similar to bottleneck adapters (Houlsby et al., 2019). We train the adapters with the standard language modeling loss $\mathcal{L}^{lm}$, where we freeze the original weights and only update the adapters. This may be analogous to finding the new embedding vector for a token with the weighted combination of original embedding vectors (Downey et al., 2023; Dobler & de Melo, 2023; Liu et al., 2024). Unlike similar works, our approach learns the weights for embedding combination. After training the adapters, we get new embeddings ${\bm{E}}^{n}\in\mathbb{R}^{|V^{n}|\times h}$ by merging the original embeddings and adapters to ${\bm{E}}^{n}={\bm{A}}{\bm{E}}^{o}$, which result in a language model with the same architecture as the original one but with a different vocabulary size.
@@ 4 Which Languages Benefit the Most from Vocabulary Adaptation?
We aim to understand “When and how should we perform vocabulary adaptation?”. More specifically, we seek insight into which languages might benefit the most from vocabulary adaptation in terms of improving overall performance or mitigating over-fragmentation.666 The non-English languages that we cover are all highly fragmented by common LLMs, and their fragmentation is simmilarly improved by our method. Therefore, our analysis focuses on performance.
@@ 6.2 Which Languages Benefit the Most from Vocabulary Adaptation?
Instead of “averaging out” the languages, we get closer look on the impact of vocabulary adaptation in a language-wise manner. Figure 4 shows the increased rate of task performance and tokenization statistics for various languages after applying different vocabulary adaptation methods. The results are shown for Latin group languages (Swahili, Indonesian, Estonian, and Haitian Creole) and Mixed group (Korean, Greek, Russian, and Bulgarian). We compare VocADT and ZeTT against the original Mistral model. The vertical axis of the increase rate is fixed to the positive range to see the benefit trends more easily and all the numbers including the negative range of the increase rate are in Appendix E. We use the FLORES development set for counting the tokens by various tokenizers where the semantic contents of every language are the same.
@@
In Appendix B, we further discuss the pronounced performance declines observed in Korean compared to Russian or Bulgarian within the same Mixed group. Despite improving fragmentation for Korean, we suspect that the linear combination assumption is insuffiencnt given the lack of representation of Korean characters in the original vocabulary.
@@ 6.3 Does Script Matter for Language grouping?
Figure 5 compares the performance and token count reduction between two non-English grouping strategies for Russian and Bulgaian: Mixed-script (ko, el, ru, bg) and Cyrillic-script (ru, bg, uk, kk) languages. For Russian, the consistent script language group performs slightly better, especially in the MT task. For Bulgarian, both grouping strategies deliver nearly identical results. Overall, the results suggest that maintaining a consistent script within a group enhances performance, though outcomes tend to beinfluenced more by the language type itself than by the grouping strategy.
@@ 8 Conclusion
We propose a simple and effective vocabulary adaptation method using a vocabulary adapter. Our approach consistently outperforms the original Mistral model after the adaptation phase across various tasks and after the following full-weight finetuning on machine translation. Furthermore, our method is on par with or more effective than strong vocabulary adaptation baselines, without relying on external embeddings or language constraints, offering a flexible and scalable solution for handling multiple languages. Our experiments cover a wide range of languages and scripts, revealing that languages with Latin scripts or severe fragmentation benefit the most. We also explored different grouping strategies, finding that maintaining consistent scripts within a group offers relatively minor benefits. Lastly, with a focus on machine translation, we confirm that vocabulary adaptation remains effective even after full weight fine-tuning, and VocADT is the most effective approach.

### v3

@@ 3.1 Vocabulary Adapter Module
We introduce the vocabulary adapter modules to find parameters of new embeddings that can replace the original embedding without changing the non-embedding part of the original model (Figure 1(a)). For simplicity, we refer to both input and output embeddings (or the LM head) collectively as embeddings. Let $V^{o}$ and $V^{n}$ be the $\bm{o}$riginal and $\bm{n}$ew vocabulary, respectively, and let $\mathcal{T}^{x}:w\rightarrow(t_{1},t_{2},\dots,t_{k})$ be a tokenizer associated with a vocabulary $V^{x}$ where $t_{j}\in V^{x},\forall j=1,\dots,k$. We put vocabulary adapter modules ${\bm{A}}\in\mathbb{R}^{|V^{n}|\times|V^{o}|}$ between the new vocabulary $V^{n}$ and the original embedding ${\bm{E}}^{o}\in\mathbb{R}^{|V^{o}|\times h}$ where $h$ is an embedding dimension, in a manner similar to bottleneck adapters (Houlsby et al., 2019). We train the adapters with the standard language modeling loss $\mathcal{L}^{lm}$, where we freeze the original weights and only update the adapters. This may be analogous to finding the new embedding vector for a token with the weighted combination of original embedding vectors (Downey et al., 2023; Dobler & de Melo, 2023; Liu et al., 2024). Unlike similar works, our approach learns the weights for embedding combination. After training the adapters, we get new embeddings ${\bm{E}}^{n}\in\mathbb{R}^{|V^{n}|\times h}$ by merging the original embeddings and adapters to ${\bm{E}}^{n}={\bm{A}}{\bm{E}}^{o}$, which results in a language model with the same architecture as the original one but with a different vocabulary size.
@@ 4 Which Languages Benefit the Most from Vocabulary Adaptation?
We aim to understand “When and how should we perform vocabulary adaptation?”. More specifically, we seek insight into which languages might benefit the most from vocabulary adaptation in terms of improving overall performance or mitigating over-fragmentation.666 The non-English languages that we cover are all highly fragmented by common LLMs, and their fragmentation is similarly improved by our method. Therefore, our analysis focuses on performance.
@@ 6.2 Which Languages Benefit the Most from Vocabulary Adaptation?
The previous section presented a macro view of performance across all languages. This section drills down to look on the impact of vocabulary adaptation in a language-wise manner. Figure 4 shows the increased rate of task performance and tokenization statistics for various languages after applying different vocabulary adaptation methods. The results are shown for Latin group languages (Swahili, Indonesian, Estonian, and Haitian Creole) and Mixed group (Korean, Greek, Russian, and Bulgarian). We compare VocADT and ZeTT against the original Mistral model. The vertical axis of the increase rate is fixed to the positive range to see the benefit trends more easily and all the numbers including the negative range of the increase rate are in Appendix E. We use the FLORES development set for counting the tokens by various tokenizers where the semantic contents of every language are the same.
@@
In Appendix B, we further discuss the pronounced performance declines observed in Korean compared to Russian or Bulgarian within the same Mixed group. Despite improving fragmentation for Korean, we suspect that the linear combination assumption is insufficient given the lack of representation of Korean characters in the original vocabulary.
@@ 6.3 Does Script Matter for Language grouping?
Figure 5 compares the performance and token count reduction between two non-English grouping strategies for Russian and Bulgarian: Mixed-script (ko, el, ru, bg) and Cyrillic-script (ru, bg, uk, kk) languages. For Russian, the consistent script language group performs slightly better, especially in the MT task. For Bulgarian, both grouping strategies deliver nearly identical results. Overall, the results suggest that maintaining a consistent script within a group enhances performance, though outcomes tend to be influenced more by the language type itself than by the grouping strategy.
@@ 8 Conclusion
We propose a simple and effective vocabulary adaptation method using a vocabulary adapter. Our approach consistently outperforms the original Mistral model after the adaptation phase across various tasks and after the following full-weight finetuning on machine translation. Furthermore, our method is on par with or more effective than strong vocabulary adaptation baselines, without relying on external embeddings or language constraints, offering a flexible and scalable solution for handling multiple languages. Our experiments cover a wide range of languages and scripts, revealing that languages with Latin scripts or severe fragmentation benefit the most. We also explored different grouping strategies, finding that maintaining consistent scripts within a group offers relatively minor benefits. Lastly, with a focus on machine translation, we confirm that vocabulary adaptation remains effective even after full-weight fine-tuning, and VocADT is the most effective approach.